# arXiv:2603.01488v1[cs.AI] 02 Mar 2026

### v1

@@ 1 Introduction
Simultaneously, the emergence of Large Language Models (LLMs) offers a breakthrough opportunity. Pre-trained on massive datasets, recent studies have leveraged LLMs to enhance exploration efficiency by generating intrinsic rewards or proposing high-level sub-goals Du et al. (2023); Tam et al. (2022); Xi et al. (2025). However, those methods mostly utilize the computational and generative ability of LLMs to facilitate explorationYao et al. (2023); Shinn et al. (2023), instead of finding the implied rules and formalizing general and reusable actions. In this paper, we want to explore whether the intrinsic world knowledge can serve as a valuable experience to assist agents in mitigating the blindness of learning from scratch, by utilizing the world knowledge and semantic information to translate natural language instructions into structured constraints and achieve policy adaptation to the new task.
@@ 3.3 Option Framework
In this framework, the learning process is divided into two levels: the high-level meta-controller learns to select optimal options, while the low-level controller optimizes policies to execute the selected options. A key limitation of existing option-based methods is that the set of options must be predefined by domain experts, relying on prior human knowledge.
@@ 4.2 Semantic Skill Generation Module
This module aims to learn policy $\pi_{o}$ considering semantics to construct a skill reuse system, similar to recent trends in using LLMs for embodied control Wang et al. (2025); Chen et al. (2025); Liang et al. (2022), which allows agents to use the skill when facing similar events. Specifically, the Semantic Skill Generation Module first explores options according to the Option Set and the current state. Then, we transmit the state changes encoding to LLMs based on hand-made prompts. LLMs then generate structured Semantic Labels $l_{o}$ for the current policy $\pi_{o}$ accordingly, which include the information of the initial and target objects. Subsequently, the module filters effective skills, based on the policy’s success rate $sr(o)$, given by interaction with environments, and stores them in the Skill Library. When utilizing Skill Library, we compare the policy and the one with the same semantic labels stored in the library based on the accuracy, and choose the more precise one as the current policy.
@@ 4.3 Constraint Adaptation Module
This module is designed to formulate adaptive rules based on intuitive linguistic instructions. Unlike traditional constrained policy optimization Yang et al. (2025). As shown in the Algorithm 1, our workflow operates by transmitting natural language as constraints to LLMs to extract abstract entities via handcrafted prompt templates, which are subsequently transformed into a set of corresponding atomic propositions, i.e., the limitation set. This set is then injected into a Reward Machine to monitor state feedback in real-time and trigger penalty signals for violations.
@@ 5.1.2 Results
  • Policy Adaptation for Extra Constraints To verify the effectiveness of our Constraint Adaptation Module in this modified environment, we designed two experimental settings: (1) Sequential Transfer in the office world B, where it learns Task1 and Task2 before Task3 (Experiment 1 in the figure); (2) Learning from Scratch in the office world B, where the agent learns Task3 directly (Experiment 2 in the figure);and (3) Learning based on previous experience, utilizing the agent have been trained in the office world A to train in the office world B. As shown in Fig.4(f), the cumulative violations in three cases eventually plateau, indicating that the agent successfully internalizes the natural language constraints. Moreover, the Sequential Transfer method results in significantly fewer total violations compared to learning from scratch, proving that reusing pre-trained skills effectively minimizes unsafe exploration.
@@ 5.2.2 Results
  • Intra-task Skill Reuse The experimental results demonstrate the framework’s capability to discover and reuse skills within a single complex task. As illustrated in Fig.5(c), the LLM-driven module generates interpretable semantic labels for learned options using a ”Predicate-Argument” structure. A critical observation is the semantic equivalence between different options. For instance, both $Opt3$ (executed without a key) and $Opt8$ (executed with a key) are annotated with the same label: act(RightLadder, MiddleLadder). This indicates that the LLM successfully abstracted the navigation policy from the low-level state variable (holding a key), allowing the agent to reuse the same ”Move from right ladder to middle ladder” skill across different task stages. This mechanism significantly reduces redundant learning of identical physical movements.

### v2

# LLM-assisted Semantic Option Discovery for Facilitating Adaptive Deep Reinforcement Learning

@@ 1 Introduction
Simultaneously, the emergence of Large Language Models (LLMs) offers a breakthrough opportunity. Pre-trained on massive datasets, recent studies have leveraged LLMs to enhance exploration efficiency by generating intrinsic rewards or proposing high-level sub-goals Du et al. (2023); Tam et al. (2022); Xi et al. (2025). However, those methods mostly utilize the computational and generative ability of LLMs to facilitate explorationShinn et al. (2023), instead of finding the implied rules and formalizing general and reusable actions. In this paper, we want to explore whether the intrinsic world knowledge can serve as a valuable experience to assist agents in mitigating the blindness of learning from scratch, by utilizing the world knowledge and semantic information to translate natural language instructions into structured constraints and achieve policy adaptation to the new task.
@@ 3.3 Option Framework
In this framework, the learning process is divided into two levels: the high-level meta-controller learns to select optimal options, while the low-level controller optimizes policies to execute the selected options.
@@ 4.2 Semantic Skill Generation Module
This module aims to learn policy $\pi_{o}$ considering semantics to construct a skill reuse system, similar to recent trends in using LLMs for embodied control, which allows agents to use the skill when facing similar events. Specifically, the Semantic Skill Generation Module first explores options according to the Option Set and the current state. Then, we transmit the state changes encoding to LLMs based on hand-made prompts. LLMs then generate structured Semantic Labels $l_{o}$ for the current policy $\pi_{o}$ accordingly, which include the information of the initial and target objects. Subsequently, the module filters effective skills, based on the policy’s success rate $sr(o)$, given by interaction with environments, and stores them in the Skill Library. When utilizing Skill Library, we compare the policy and the one with the same semantic labels stored in the library based on the accuracy, and choose the more precise one as the current policy.
@@ 4.3 Constraint Adaptation Module
This module is designed to formulate adaptive rules based on intuitive linguistic instructions. Unlike traditional constrained policy optimization. As shown in the Algorithm 1, our workflow operates by transmitting natural language as constraints to LLMs to extract abstract entities via handcrafted prompt templates, which are subsequently transformed into a set of corresponding atomic propositions, i.e., the limitation set. This set is then injected into a Reward Machine to monitor state feedback in real-time and trigger penalty signals for violations.
@@ 5.1.2 Results
  • Policy Adaptation for Extra Constraints To verify the effectiveness of our Constraint Adaptation Module in this modified environment, we designed three experimental settings: (1) Sequential Transfer in the office world B, where it learns Task1 and Task2 before Task3 (Experiment 1 in the figure); (2) Learning from Scratch in the office world B, where the agent learns Task3 directly (Experiment 2 in the figure); and (3) Learning based on previous experience, utilizing the agent have been trained in the office world A to train in the office world B. As shown in Fig.4(f), the cumulative violations in three cases eventually plateau, indicating that the agent successfully internalizes the natural language constraints. Moreover, the Sequential Transfer method results in significantly fewer total violations compared to learning from scratch, proving that reusing pre-trained skills effectively minimizes unsafe exploration.
@@ 5.2.2 Results
  • Intra-task Skill Reuse The experimental results demonstrate the framework’s capability to discover and reuse skills within a single complex task. As illustrated in Fig.5(c), the LLM-driven module generates interpretable semantic labels for learned options using a ”Predicate-Argument” structure. A critical observation is the semantic equivalence between different options. For instance, both $act3$ (executed without a key) and $act8$ (executed with a key) are annotated with the same label: act(RightLadder, MiddleLadder). This indicates that the LLM successfully abstracted the navigation policy from the low-level state variable (holding a key), allowing the agent to reuse the same ”Move from right ladder to middle ladder” skill across different task stages. This mechanism significantly reduces redundant learning of identical physical movements.